find a value for K that will make 4x^2 + 6.4x + K a perfect square. Describe the procedure that you used which require lgebra[that is, not trial and error]
Answer by stanbon(75887)

find a value for K that will make 4x^2 + 6.4x + K a perfect square. Describe --the procedure
--------------
A perfect square would look like this: a^2 + 2ab + b^2
In your problem a^2 = 4x^2, so a = 2x
In your problem 2ab = 6.4x
So, 2(2x)b = 6.4x
So, 4b = 6.4
So, b = 1.6
-------------------
In your equation b^2 = K
So, K = (1.6)^2 = 2.56
============================
Cheers,
Stan H.